RektbyFaith:

2 Pre Calc problems I need help?

RektbyFaith:

1 attachment
Shadow:

@Vocaloid

Shadow:

@Hero

ThewaffleBRO:

@TheWaaffleBRO

ThewaffleBRO:

oh wait thats me

RektbyFaith:

lol

RektbyFaith:

How many people does it take to solve 2 pre calc problems? Let's find out folks!

ThewaffleBRO:

folks

ThewaffleBRO:

did i go back in time

RektbyFaith:

it's possible, I mean this math is ancient enough to be your age :thinking:

Ultrilliam:

@Angle

Vocaloid:

yeah that's correct (sorry that took so long, had to re-teach myself this)

RektbyFaith:

What do you think @Vocaloid ?

1 attachment
ThewaffleBRO:

i think

RektbyFaith:

if someone says a lot of numbers they get thrown into a volcano.

Ultrilliam:

A lot of numbers or 27582375890347892347683476893476843763486734897623489672348

RektbyFaith:

It looks like it's D honestly.

Hero:

Looks right @RektbyFaith Demoivre's Formula: \(\cos(x) + i\sin(x))^n = \cos(nx) + i\sin(nx)\)

RektbyFaith:

yeah but @Hero look at this one

1 attachment
RektbyFaith:

I think D?

Angle:

m = a[cos(x) + i sin(x)] n = b[cos(y) + i sin(y)] then m*n = (a*b)[cos(x+y) + i sin(x+y)]

Angle:

D is correct :)

RektbyFaith:

oooo yay, last one im so sorry lol. I didn't notice it

RektbyFaith:

1 attachment
RektbyFaith:

C and A?

Hero:

@RektbyFaith you're posting too many problems. I thought you were working on the first one you posted.

RektbyFaith:

yeah sorry lol @Hero I'll be out of your hair :)

Angle:

Since the question poster can only give one medal per post, we prefer if people only ask question per question thread. Feel free to close this question thread and ask a new one :)

RektbyFaith:

Okay np :)